OpenStudy (anonymous):

Which pair of numbers is relatively prime? A. 10 and 35 B. 10 and 27 C. 10 and 70 D. 10 and 48

OpenStudy (anonymous):

I don't know what relatively prime mean

Directrix (directrix):

Two numbers are relatively prime if they have no common factors except for 1. So, which pair do you think fits that description? @hannahekelman

OpenStudy (anonymous):

I think its c then

Directrix (directrix):

These have 5 in common --> C. 10 and 70 Try again.

Directrix (directrix):

Two numbers are "relatively prime" if they have no common factors other than 1 (in other words you cannot evenly divide both by some common value) Example: 7 and 20 are relatively prime (no common factor), but 6 and 20 are not relatively prime because you can divide both by 2 (2 is a common factor).

Directrix (directrix):

Go through your options: Do these two numbers have a common factor: 10 and 35 ? @hannahekelman

Directrix (directrix):

Look at these two: 10 and 27 10 = 5*2 27 = 3*3*3 Do 10 and 27 have a common factor shown above? @hannahekelman